# A Theory of Recommendations

### v2

@@ 3 Model
We begin by introducing the assumptions of our benchmark model in Section 3.1. Then, in Section 3.2, we discuss and motivate these assumptions.
3.1 Assumptions
@@
We consider a consumer’s choice of a product from a set of alternatives $N\in\mathbb{N}$ with $N>1$. We assume that each product is fully characterized by a two-dimensional quality vector, where $Q_{d}=1$ and $Q_{d}=0$ capture high and low quality, respectively, in dimension $d\in\{1,2\}$. Accordingly, four different product versions exist. We assume that the products are experience goods. Thus, the consumer cannot observe product quality before consumption, so the alternatives are identical for him ex ante. However, he knows the prior distribution of the quality vectors, referred to as quality distribution for brevity. Consumers have heterogeneous preferences fully characterized by their type $i\in[-1/2,1/2]$. A consumer of type $i$ receives a payoff of
@@
We assume that $\mathbf{q}>0$ unless mentioned otherwise.
@@
We assume that one of the products comes with a recommendation. The receiver has to choose between this product and an alternative without a recommendation. We distinguish between buy recommendations ($r=B$) and don’t-buy recommendations ($r=D$). Moreover, $r=0$ corresponds to the case without recommendation. In the benchmark model, we assume that the senders (she) and receivers (he) of recommendations are both randomly drawn from $F$.
@@
We assume that receivers are Bayesian, updating beliefs about product quality after having obtained a recommendation while taking into account the uncertainty about its informational content. We define $\pi^{B}(R)$ and $\pi^{D}(R)$ as the probability that a randomly chosen sender gives a buy or don’t-buy recommendation, respectively. Clearly,
@@
$\pi^{D}(R)$ $:=q_{1}(1-\phi_{1}(R))+q_{2}(1-\phi_{2}(R))+q_{L},$
where $\pi^{D}(R)=1-\pi^{B}(R)$, and $\phi_{1}(R):=1-F(R-1/2)$ and $\phi_{2}(R):=F(1/2-R)$ denote the probability of a buy recommendation conditional on the product version being $(1,0)$ or $(0,1)$, respectively. In words, the probability of receiving a buy recommendation from a randomly chosen sender is given by the sum of the probability that a product is objectively good and the probability that it is subjectively good enough from a randomly chosen sender’s perspective. Thus, accepting a buy recommendation the receiver’s posterior $\mathbf{p}^{B}(R):=(p_{H}^{B}(R),p_{1}^{B}(R),p_{2}^{B}(R),p_{L}^{B}(R))$ becomes
In the above expressions $\pi^{B}(R)$, $p^{B}(R)$, etc., we will typically drop the dependence on $R$ to simplify notation. Finally, we use the following summary terminology:
@@
3.2 Discussion of the Framework
Our theory of recommendations incorporates three key features: heterogeneity of products and preferences, coarseness of recommendations, and asymmetry of stakes between senders and receivers. We now discuss how our model addresses these and further points.
Heterogeneity of Products and Preferences
Without controversial products or preference heterogeneity, the analysis would become trivial. The case without controversial products would correspond to $q_{1}=q_{2}=0$. Then, for arbitrary $R\in(0,1),$ buy recommendations would reveal that the product is good, while don’t-buy recommendations would reveal that it is bad. Thus, there would be no friction in the recommendation system and no role for design considerations. With controversial products but absent preference heterogeneity, all consumers would have the same type $i^{*}$. Assuming w.l.o.g. that $i^{*}\geq 0$, senders would send a buy recommendation (i) only after consuming a good product if $R>\frac{1}{2}+i^{*}$, (ii) after consuming a good product or a controversial product of version $(1,0)$ if $\frac{1}{2}+i^{*}\geq R>\frac{1}{2}-i^{*}$, and (iii) after consuming the good product or any of the controversial products if $\frac{1}{2}-i^{*}\geq R$. The first-best could be induced by the appropriate choice of the recommendation threshold, taking into account how the payoffs from controversial products compare to the expected payoff from the outside option, In sum, a meaningful analysis requires controversial products as well as preference heterogeneity.
Coarse Information Transmission
With a single threshold above and below which senders give buy and don’t-buy recommendations, respectively, it is impossible to always reveal the product version. Thus, as desired, recommendations provide coarse information. If, in contrast, the sender could give four different ratings depending on the consumption experience, full revelation would be possible in our parsimonious model with only four product versions. In reality, recommendations can convey more content than whether a buyer liked a product or not, for instance, by including verbal descriptions. However, given the typical complexity of products, even relatively detailed recommendations will usually only transmit coarse information.
Asymmetry of Stakes
To capture the sender’s limited stakes in the process, we assume that she behaves truthfully, mechanically reporting whether her consumption experience was good enough or not. More sophisticated and potentially strategic recommendations seem at odds with the lack of incentives to invest in this reporting activity, which results in the underprovision of recommendations (see the literature review).
Recommendation Threshold
Our preferred way of thinking about the recommendation threshold is as a design parameter. To motivate this, we rely on Decker (2022) who argues that Airbnb advises its customers to give a five-star rating to every host that provides reasonably good services, thus reserving lower ratings exclusively to very bad hosts. [^10] In our setting, this approach corresponds to the use of a coarse recommendation system, with the vast majority of hosts receiving a buy recommendation and a small subset receiving a don’t-buy recommendation. The recommendation threshold would then be close to $R=0$ so that products with a don’t-buy recommendation are typically objectively bad. In this spirit, we think of the threshold as reflecting the guidance of a platform such as Airbnb or as a social norm that is being actively shaped.
Other Recommendation Systems
Assuming that senders mechanically provide a binary recommendation is adequate in view of their limited willingness to invest more time into giving a recommendation. Nonetheless, one could imagine more sophisticated recommendation systems. For instance, the sender might recommend a product when it is either good or when it is controversial but preferred by most of the population. This, however, presupposes a sender with detailed information about the receiver’s decision environment. Instead, the sender could give a buy recommendation when the average payoff she expects receivers to obtain from consuming the product in question is higher than for the outside option. In that case, the expected value of the outside option would determine the recommendation threshold, which would then no longer be a design parameter.
Product Ranking
When $q_{H}>0$ and $q_{L}>0$, all types agree on which product version is the best and which is the worst because their utility is fully determined by the qualities in each dimension and is increasing in each of them. In other environments, no products are objectively the best or the worst. For instance, there would be no such products in a modified version of our setting where $q_{H}=0$ and $q_{L}=0$, as the valuations for $(1,0)$ and $(0,1)$ are perfectly negatively correlated. We nonetheless keep the specific feature of the model because we believe it covers a wide range of plausible scenarios.
4 Optimal Receiver Behavior
In this section, we ask under which circumstances a receiver will accept all recommendations, depending on his type $i$ and the recommendation system $\mathcal{R}$, as introduced in Definition 1. In Section 4.1, we discuss through which channels recommendations affect beliefs, distinguishing between the objective and subjective content of the recommendations. Section 4.2 then characterizes the optimal behavior of the receiver by identifying thresholds which divide the type set into receivers that accept all recommendations and those that never do.
4.1 The Effects of Recommendations on Beliefs
Recommendations contain coarse information about the product under consideration, leading the receiver to update his beliefs about its quality. We first focus on buy recommendations, the total effect of which can be decomposed as follows: [^11] In Appendix A.1, we provide a more formal analysis, decomposing the difference between posteriors and priors into three corresponding components.
  1. (I)
    The probability of a bad product falls to zero.
  2. (II)
    A good product becomes more likely relative to the controversial products.
  3. (III)
    The odds between the controversial products change.
The first two effects in this belief-updating decomposition are objectively positive: They correspond to increases of the expected payoff for every receiver type, albeit of different sizes. By contrast, the third effect is fully type-dependent and thus subjective, resulting in an increase in the payoff for some receivers and a decrease for others.
The effects of don’t-buy recommendations are analogous to those of buy recommendations: First, the probability of a good product falls to zero; second, bad products become more likely relative to controversial products, and third, the odds between the controversial products change. Thus, again, the recommendation has two objective effects (but these are both bad now) and a subjective effect (which can be good or bad, depending on the receiver’s type).
4.2 Characterizing Optimal Receiver Behavior
@@
We are now in a position to understand whether it is optimal for the receiver to accept recommendations. To this end, note that, after receiving a recommendation, the receiver can either buy the recommended product or not buy it. Denote by $U^{B}(i)$ and $U^{D}(i)$ player $i$’s expected payoff of buying a product with a buy and don’t-buy recommendation, respectively, and by $U^{0}(i)$ the expected payoff of not buying, i.e., of buying the alternative product with no recommendation. We have
@@
Now, the difference between $U^{r}(i)$ and $U^{0}(i)$ determines the behavior of an expected-payoff maximizing receiver as follows.
Definition 2.
We say that a receiver $i$
@@
    accepts a buy recommendation if $U^{B}(i)\geq U^{0}(i)$; otherwise we say that he does not accept a buy recommendation;
@@
    accepts a don’t-buy recommendation if $U^{0}(i)\geq U^{D}(i)$; otherwise we say that he does not accept a don’t-buy recommendation
@@
    accepts all recommendations if he accepts both buy and don’t-buy recommendations;
  4. (iv)
    never accepts recommendations if he does not accept either buy and don’t-buy recommendations.
Hence, a receiver who accepts all recommendations “does as he is told”, whereas a receiver who never accepts recommendations is a “contrarian” who does the opposite of what he is told. [^12] The use of the term aligns with Drehmann et al. (2005), who document contrarian behavior, i.e., investors going against the behavior of the predecessors, in an financial-markets experiment. Note that the above definition does not cover all conceivable types of behavior, as a receiver $i$ could accept buy recommendations and not accept don’t-buy recommendations, which is tantamount to always buying the recommended product and thus ignoring the content of the recommendations. Analogously, a receiver $i$ could ignore the content of the recommendations and never buy the recommended product. We will return to these cases shortly.
The condition from Definition 2(i) for receiver $i$ to accept a buy recommendation can equivalently be written as
For more compact formulations, we introduce the following terminology.
@@
    either accepts all recommendations or never accepts recommendations;
@@
    accepts all recommendations if and only if $\Delta_{O}^{B}\geq i\Delta_{S}^{B}$.
The first part of the result states that Definition 2 indeed captures all relevant forms of behavior. That is, no type of receiver will ignore the informational content of recommendations by always or never buying the recommended product. Intuitively, a receiver accepts all recommendations if he thinks that they have sufficient objective content and/or that his preferences are sufficiently aligned with the sender’s. Conversely, a receiver never accepts recommendations if he thinks that they do not have sufficient objective content and/or that his preferences are not sufficiently aligned with the sender’s. Whether the recommendation is positive or negative does not matter for this assessment. Rather, as laid out in the decomposition in Section 4.1, the objective effect $\Delta_{O}^{B}$ of a buy recommendation is positive, as the latter makes a good product more likely relative to the controversial products and rules out a bad one. By contrast, the sign of the subjective effect $\Delta_{S}^{B}$ can be positive or negative. The second part of the result shows which types accept all recommendations, which happens unless the subjective effect is sufficiently large and the receiver’s preferences are not sufficiently well aligned with those of the rest of the population. We emphasize that receivers who never accept recommendations still make use of them, by doing the opposite of what they are told.
Proposition 1 tells us that we can decompose the population into at most two different groups: those who accept all and those who never accept recommendations. Denoting the type who is indifferent between always accepting and never accepting recommendations as
@@
    If $|\Delta_{S}^{B}|\leq 2\Delta_{O}^{B}$, then all $i\in[-1/2,1/2]$ accept all recommendations.
@@
    If $\Delta_{S}^{B}<-2\Delta_{O}^{B}$, then all $i\geq\tilde{i}$ accept all recommendations.
@@
    If $\Delta_{S}^{B}>2\Delta_{O}^{B}$, then all $i\leq\tilde{i}$ accept all recommendations.
The intuition for Corollary 1, the proof of which is omitted, is straightforward. Without loss, consider a buy recommendation. For $\Delta_{S}^{B}=0$, the recommendation does not shift probability between the controversial product versions. Accordingly, accepting the recommendation is optimal for all players because of its positive objective effect $\Delta_{O}^{B}\geq 0$. For $\Delta_{S}^{B}\neq 0$, the conclusion is the same as long as $|\Delta_{S}^{B}|$ remains small enough—as required by (i). In that case, even the receivers who are most adversely affected by the subjective content of the recommendation (those at $i=1/2$ or $i=-1/2$) are convinced by its relatively objective character. As $|\Delta_{S}^{B}|$ increases, the role of the probability shift becomes apparent. As types $i>0$ prefer (1,0) over (0,1), a shift of the probability from the latter to the former induced by a recommendation ($\Delta_{S}^{B}<0$) increases such receivers’ expected payoff, reinforcing the objective effect. By contrast, such a shift reduces the expected payoff of types $i<0$. The condition in (ii) always holds for $i>0$, as the objective and subjective effects reinforce each other. It also holds for $i<0$ as long as $|i|<|\tilde{i}|$ so that the adverse subjective effect does not dominate the objective effect. Similarly, when the recommendation shifts the probability from (1,0) to (0,1) (when $\Delta_{S}^{B}>0$), all types $i<0$ accept. By contrast, types $i>0$ experience a reduction in their value, which can outweigh the positive objective effect. The condition in (iii) makes sure that the latter case does not arise. [^15] The effect on the expected payoffs in cases (ii) and (iii) is similar to the effect of a demand rotation as discussed in Johnson and Myatt (2006): For the types who accept the recommendation, the expected payoffs increase; for those who do not, they fall.
Corollary 1 may potentially be useful to draw inferences about the relative importance of the objective and subjective effects of recommendations: If an observer sees a large share of receivers accepting recommendations, this can be interpreted as evidence for a relatively high objective effect of the recommendation.
Finally, we analyze receiver behavior for extreme recommendation thresholds.
@@
Suppose $R\rightarrow 1$ or $R\rightarrow 0$. Then, all types accept all recommendations irrespective of the product distribution.
In the limits, both types of recommendations essentially contain only objective information. For instance, when $R\rightarrow 1$, buy recommendations require that the product is objectively good, while don’t-buy recommendations require that it is not. Either way, the recommendation does not reveal anything about the relative probabilities of the two controversial product versions. Instead, it provides objective information that is equally valuable for all types, irrespective of their preferences for controversial products. Thus, all types accept all recommendations.
5 The Value of the Recommendation System
We now deal with the value of recommendation systems. In Section 5.1, we provide a characterization result. Section 5.2 asks how to use the recommendation threshold to influence the value. Finally, we investigate the effects of the preference and quality distributions on the value in Section 5.3.
5.1 Characterizing the Value of the Recommendation System
Before defining the value of a recommendation system, we describe the expected payoffs of a fixed type $i$ receiver, depending on whether he has access to a recommendation system and whether he accepts its recommendations. Without a recommendation system, all products are identical ex ante, with an expected payoff based on the prior distribution of product qualities. Hence, the expected payoff of a receiver $i$ is
@@ 5.1 Characterizing the Value of the Recommendation System
Now suppose one of the products comes with a buy or don’t-buy recommendation, so the receiver has to update his belief about the quality distribution. A receiver who accepts all recommendations chooses this product (rather than one of the alternatives) only if it comes with a buy recommendation. With a recommendation system, the expected payoff of a receiver $i$ who accepts all recommendations is thus
@@
The first term is the probability of receiving a buy recommendation multiplied by the expected payoff from accepting it (and thus buying the recommended product). The second term is the probability of a don’t-buy recommendation multiplied by the expected payoff from buying the product without any recommendation, that is, based on priors. [^16] The notation $V^{A}(i)$ emphasizes the dependence of this expected payoff on the type of the receiver. Clearly, $V_{A}$ also depends on $R$ and on the decision environment; similarly for the expressions below.
A receiver who never accepts recommendations will buy the recommended product if it comes with a don’t-buy recommendation. Otherwise, he will buy one of the other products. Hence, the expected payoff of such a receiver is
@@
A buy recommendation does not create value for receivers who do not accept it, as such types choose the alternative for which no recommendation is available. Instead, the don’t-buy recommendation, which is given with probability $(1-\pi^{B})$, does influence the behavior of the types who do not accept it: Instead of buying an alternative based on prior probabilities, updating leads them to choose the product for which they received the explicit recommendation not to buy.
We now define the value of a recommendation system. [^17] This definition is in line with the definition of a valuable signal in Kamenica and Gentzkow (2011).
@@
The simple structure of $V(\mathcal{R})$ for $|\Delta_{S}^{B}|\leq 2\Delta_{O}^{B}$ in (i) results because, by Corollary 1, all types accept a recommendation. The difference between the expressions for $\Delta_{S}^{B}<-2\Delta_{O}^{B}$ in (ii) and $\Delta_{S}^{B}>2\Delta_{O}^{B}$ in (iii) arises because only types $i>\tilde{i}$ accept the recommendation in the former case, whereas only types $i<\tilde{i}$ do so in the latter.
We now express the value $V(\mathcal{R})$ in terms of its objective and subjective content.
@@
The value $V(\mathcal{R})$ of the recommendation system $\mathcal{R}$ is:
@@
Intuitively, when $|\Delta_{S}^{B}|\leq 2\Delta_{O}^{B}$, Corollary 1 implies that all receivers accept all recommendations as the objective effect dominates. Hence, by equation (6), the recommendation system creates value only in case of a buy recommendation, as a don’t-buy recommendation leads to the same expected payoff as without a recommendation system. The positive payoff effect of the recommendation reflects the type-independent objective part $\Delta_{O}^{B}$ and the subjective part $\Delta_{S}^{B}$, which is weighted by the average type. Finally, all of that only materializes with the probability of a buy recommendation, $\pi_{B}$.
As the absolute value of $\Delta_{S}^{B}$ increases, the analysis becomes more complex because some types never accept recommendations. We consider the case $\Delta_{S}^{B}>2\Delta_{O}^{B}$ and reproduce the expression for $V(\mathcal{R})$ as follows: [^18] The argument for $\Delta_{S}^{B}<-2\Delta_{O}^{B}$ is analogous.
$\underbrace{\pi^{B}F(\tilde{i})\left[\Delta_{O}^{B}-\Delta_{S}^{B}\mathbb{E}[i\mid i\leq\tilde{i}]\right]}_{\text{types accept all recommendations}}+\underbrace{(1-\pi^{B})(1-F(\tilde{i}))\left[\Delta_{O}^{D}-\Delta_{S}^{D}\mathbb{E}[i\mid i\geq\tilde{i}]\right]}_{\text{types never accept recommendations}}.$ (8)
In this case, the buy recommendation shifts probability from $(1,0)$ to $(0,1)$, which is unfavorable for receivers $i>0$, reducing their gains from accepting and possibly making them negative. Thus, for $\Delta_{S}^{B}>2\Delta_{O}^{B}$, only types below $\tilde{i}$ (a fraction $F(\tilde{i})$ of the receivers) accept all recommendations, which comes with probability $\pi^{B}$. In addition to the objective value of the recommendation ($\Delta_{O}^{B}$), these receivers obtain a subjective value. In (8), the subjective value is reflected in $-\Delta_{S}^{B}\mathbb{E}[i\mid i\leq\tilde{i}]$, the average effect of the shift in the relative probabilities of $(1,0)$ and $(0,1)$ on the value of accepting the recommendation for these buyers. The sign of this term depends on the properties of $F$ captured in $\mathbb{E}[i\mid i\leq\tilde{i}]$. It is positive if the negative contribution of types $i<0$ dominates the positive contribution of types $i>0$. The second component in (8), corresponding to the types who never accept recommendations, is arguably more surprising. Given their lack of alignment with the rest of the population, these types treat the recommendation not to buy as good news about the product and thus do not accept the recommendation. [^19] In detail, a don’t-buy recommendation (which happens with probability $1-\pi^{B}$) carries an adverse objective content (summarized by $\Delta_{O}^{D}$ in the term in squared brackets). However, for receivers with $i>0$ it also carries the good news that the product is more likely to be of the preferred version $(1,0)$ rather than $(0,1)$. Those with $i>\tilde{i}(>0)$ then buy the product despite the adverse objective content, valuing the beneficial subjective effect with $-\Delta_{S}^{D}\mathbb{E}[i\mid i\geq\tilde{i}]>0$ on average. Such receivers exist only if $\Delta_{S}^{B}$ is sufficiently large (and hence $\Delta_{S}^{D}$ is sufficiently negative relative to $\Delta_{O}^{D}$). Nonetheless, its informational content creates value for them.
Going forward, we shall often invoke the following symmetry assumption on the preference distribution.

### v3

@@ 2 Model
We consider a consumer’s choice between two products. We assume that each product is fully characterized by a two-dimensional quality vector, where $Q_{d}=1$ and $Q_{d}=0$ capture high and low quality, respectively, in dimension $d\in\{1,2\}$. Accordingly, four different product versions exist. The consumer cannot observe product quality before consumption, so the alternatives are identical for him ex ante. However, he knows the prior distribution of the quality vectors, referred to as quality distribution for brevity. Consumers have heterogeneous preferences fully characterized by their type $i\in[-1/2,1/2]$. A consumer of type $i$ receives a payoff of
@@
We assume that $\mathbf{q}>0$ unless mentioned otherwise. We sometimes use the following self-explanatory terminology:
Definition 1.
@@
We assume that one of the products comes with a recommendation. The receiver has to choose between this product and the alternative without a recommendation. We distinguish between buy recommendations ($r=B$) and don’t-buy recommendations ($r=D$). Moreover, $r=0$ corresponds to the case without recommendation. In the benchmark model, we assume that the senders (she) and receivers (he) of recommendations are both randomly drawn from $F$.
@@
We assume that receivers are Bayesian, updating beliefs about product quality after having obtained a recommendation while taking into account the uncertainty about its informational content. The probability that a randomly chosen sender gives a buy or don’t-buy recommendation, respectively is given as
@@
$\pi^{D}(R)$ $:=1-\pi^{B}(R)=q_{1}(1-\phi_{1}(R))+q_{2}(1-\phi_{2}(R))+q_{L}.$
Here $\phi_{1}(R):=1-F(R-1/2)$ and $\phi_{2}(R):=F(1/2-R)$ denote the probability of a buy recommendation conditional on the product version being $(1,0)$ or $(0,1)$, respectively. In words, the probability of receiving a buy recommendation from a randomly chosen sender is given by the sum of the probability that a product is objectively good and the probability that it is subjectively good enough from a randomly chosen sender’s perspective. In the appendix, we present the receiver’s posterior $\mathbf{p}^{r}(R):=(p_{H}^{r}(R),p_{1}^{r}(R),p_{2}^{r}(R),p_{L}^{r}(R))$ after a buy recommendation $r=B$ or a don’t-buy recommendation $r=D$ [^10] To ease notation, we will typically drop the dependence of these posteriors and other variables on $R$.
Finally, we use the following summary terminology:
@@
3 Preliminaries: Receiver Behavior, Value and Demand
In this section, we begin by asking under which circumstances a receiver will accept recommendations, depending on his type $i$ and the recommendation system $\mathcal{R}$. Then, we will use this result to characterize the value of a recommendation system and determine the demand induced by the recommendation system.
3.1 Receiver Behavior
@@
Denote by $U^{B}(i)$ and $U^{D}(i)$ player $i$’s expected payoff of buying a product with a buy and don’t-buy recommendation, respectively, and by $U^{0}(i)$ the expected payoff of not buying, i.e., of buying the alternative product with no recommendation. We have
@@
Receiver $i$ will accept a buy recommendation if $U^{B}(i)\geq U^{0}(i)$ or equivalently
Intuitively, a buy recommendation conveys the following information to the receiver:
@@
    The probability of a bad product has fallen to zero.
@@
    The likelihood of a good product relative to the controversial products has increased.
@@
    The odds between the controversial products has changed.
Similarly, the receiver will accept a don’t-buy recommendation if $U^{0}(i)\geq U^{D}(i)$, and the effects of these recommendations are analogous. For more compact formulations, we introduce the following terminology.
@@
For the unbiased receiver $i=0$, equation (4) simplifies to $\Delta_{O}^{B}\geq 0$. As this receiver does not care about the shift in probabilities between the two controversial versions of the products, $\Delta_{O}^{B}\geq 0$ reflects the two objective effects (I) and (II). For $i\neq 0$, the interpretation of $\Delta_{O}^{B}\geq 0$ is more nuanced. If $q_{1}=q_{2}$, the expected payoff of a controversial product is type-independent, ensuring that $\Delta_{O}^{B}$ fully captures the objective benefits of a buy recommendation. If $q_{1}\neq q_{2}$, the size of these benefits depends on the receiver’s type. [^12] Intuitively, if one controversial version (e.g., $(1,0)$) is ex ante more likely, learning that a product is controversial rather than bad is more valuable for those who prefer $(1,0)$ over $(0,1)$, and vice versa for the opposite preference. Similarly, learning that a product is good rather than controversial is more valuable for those who favor the less likely version. Nonetheless, the sign of $\Delta_{O}^{B}$ remains a reliable indicator of the objective effect, as all types agree that (I) and (II) are beneficial. We can now state our first result.
@@
    accepts buy recommendations if and only if he accepts don’t-buy recommendations;
@@
    accepts (buy and don’t-buy) recommendations if and only if $\Delta_{O}^{B}\geq i\Delta_{S}^{B}$.
Intuitively, a receiver accepts recommendations if he thinks that they have sufficient objective content and/or that his preferences are sufficiently aligned with the sender’s. Conversely, a receiver does not accept recommendations if he thinks that they do not have sufficient objective content and/or that his preferences are not sufficiently aligned with the sender’s. Whether the recommendation is positive or negative does not matter for this assessment. Denoting the type who is indifferent between accepting and not accepting recommendations as
@@
    If $|\Delta_{S}^{B}|\leq 2\Delta_{O}^{B}$, then all $i\in[-1/2,1/2]$ accept recommendations.
@@
    If $\Delta_{S}^{B}<-2\Delta_{O}^{B}$, then all $i\geq\tilde{i}$ accept recommendations.
@@
    If $\Delta_{S}^{B}>2\Delta_{O}^{B}$, then all $i\leq\tilde{i}$ accept recommendations.
The intuition for Corollary 1, the proof of which is omitted, is straightforward. As long as the subjective content of the recommendation is small (case (i)), even the most extreme receivers satisfy the cut-off condition of Proposition 1(ii). As the subjective content increases, consumers who are not sufficiently aligned with the population at large will reject recommendations (cases (ii) and (iii)).
The interplay between objective and subjective effect leads to a stark result for extreme recommendation thresholds:
@@
Suppose $R\rightarrow 1$ or $R\rightarrow 0$. Then, all types accept recommendations irrespective of the product distribution.
In the limits, recommendations contain only objective information. As $R\to 1$, a buy recommendation guarantees the product is good, while a don’t-buy recommendation implies it is not. Since no information is revealed about the relative likelihood of controversial products, all types value the recommendation equally and accept it.
Next, we discuss how the analysis would become trivial in the absence of either controversial products or preference heterogeneity. If there were no controversial products ($q_{1}=q_{2}=0$), recommendations would simply reveal whether the product is good or bad. In this case, recommendations would contain only objective information, eliminating any friction in the system. If there was no preference heterogeneity, the sender’s recommendations would convey objective information about whether the product is good, bad, or controversial. Without heterogeneity in receiver preferences, the subjective effect of a recommendation—how it shifts beliefs between different controversial product versions—would be irrelevant, as all receivers would interpret the information in the same way. As a result, the sender could fully align recommendations with the receivers’ interests by adjusting the recommendation threshold, effectively eliminating any tension. Hence, it is the interplay between controversial products or preference heterogeneity that makes the design of the recommendation system nontrivial and interesting.
3.2 Value of the Recommendation System
Having determined receiver behavior, we can now characterize the value of the recommendation system. Without a recommendation system, all products are identical ex ante, with an expected payoff based on the prior distribution of product qualities. Hence, the expected payoff of a receiver $i$ is
@@ 3.2 Value of the Recommendation System
Now suppose one of the products comes with a buy or don’t-buy recommendation, so the receiver has to update his belief about the quality distribution. A receiver who accepts recommendations chooses this product (rather than one of the alternatives) only if it comes with a buy recommendation. With a recommendation system, the expected payoff of a receiver $i$ who accepts recommendations is thus
@@
The first term on the right-hand side comes from buy recommendations, the second one from don’t-buy recommendations.
A receiver who does not accept recommendations will buy the recommended product if it comes with a don’t-buy recommendation. Otherwise, he will buy one of the other products. Hence, the expected payoff of such a receiver is
@@
A buy recommendation does not create value for receivers who do not accept it, as such types choose the alternative for which no recommendation is available. Instead, the don’t-buy recommendation, which is given with probability $(1-\pi^{B})$, does influence the behavior of the types who do not accept it: Rather than buying an alternative based on prior probabilities, updating leads them to choose the product for which they received the explicit recommendation not to buy.
We now define the value of a recommendation system in terms of the expected welfare increase it creates for receivers. [^13] This definition is in line with the definition of a valuable signal in Kamenica and Gentzkow (2011).
@@
Importantly, the definition captures the idea that the increase in expected utility is measured before drawing the sender-receiver pair. Part (i) relies on Corollary 1, which implies that all receivers accept recommendations when $|\Delta_{S}^{B}|\leq 2\Delta_{O}^{B}$, as the objective effect dominates. Hence, by equation (6), the recommendation system creates value only in case of a buy recommendation, as a don’t-buy recommendation leads to the same expected payoff as without a recommendation system. Parts (ii) and (iii) refer to cases where the subjective content of the messages matters more. Then, not all receivers accept the recommendation, and the system may create value for both receivers who accept recommendations as well as those who do not. We restate the value $V(\mathcal{R})$ to better understand how it reflects the role of the objective and subjective content of recommendations:
@@
The value $V(R)$ of the recommendation system $\mathcal{R}$ is: [^14] As we focus on $R$ as the design parameter, we often write $V(R)$ instead of $V(\mathcal{R})$.
@@
The positive payoff effect of the recommendation in case (i) reflects the type-independent objective part $\Delta_{O}^{B}$ and the subjective part $\Delta_{S}^{B}$, which is weighted by the average type. As the subjective component $\Delta_{S}^{B}$ becomes stronger in absolute terms, not all receiver types will follow the recommendation anymore, and the analysis becomes more nuanced. To illustrate this, consider case (iii) where $\Delta_{S}^{B}>2\Delta_{O}^{B}$. In this case, a buy recommendation shifts beliefs toward the controversial product version that is less preferred by receivers $i>0$, which implies that they may decide not to follow the recommendation. Only those whose preferences are sufficiently aligned with the sender’s—specifically, those with types below a certain threshold $\tilde{i}$—still choose to buy. These types represent a fraction $F(\tilde{i})$ of the population. For them, the recommendation carries both objective value (because the product is likely better than the outside option) and subjective value (because the controversial version it signals is relatively well aligned with their preferences). This subjective value is captured by the average type among those who follow the recommendation, $\mathbb{E}[i\mid i\leq\tilde{i}]$.
Even those types who reject the buy recommendation may benefit. Types above $\tilde{i}$, who choose not to follow the recommendation, infer that the product is likely of the version they personally dislike—yet this very inference tells them something useful. When these same types receive a don’t-buy recommendation, they update in favor of the controversial product version they prefer, and may end up buying despite, or, rather, because of the negative signal. So, while the objective signal says “don’t buy”, the subjective shift in beliefs may outweigh it for some types. As a result, even rejected recommendations can create value, precisely because they help those whose preferences diverge from the majority make more informed choices.
3.3 Demand for the Recommended Product
A platform does not necessarily design a recommendation system to maximize the value it creates for consumers. Rather, the objective may be to maximize profits, which in some circumstances amounts to maximizing demand. [^15] For instance, assuming that a platform has committed to a fixed transaction fee, the profit-maximizing platform will want to to adjust its recommendation system to maximize demand. Whether a receiver buys the recommended product depends on the content of the recommendation and whether the receiver accepts or rejects the recommendation. Therefore, demand may comprise receivers who accept and receivers who reject recommendations. Using Corollary 1, the demand for the recommended product $D(R)$ can be expressed analogously to the value of the recommendation system in Proposition 3 as
$D(R)=\begin{cases}\pi_{B}(R)&\text{if }|\Delta_{S}^{B}|\leq 2\Delta_{O}^{B}\\ \pi_{B}(R)F(\tilde{i}(R))+(1-\pi_{B}(R))(1-F(\tilde{i}(R)))&\text{if }\Delta_{S}^{B}>2\Delta_{O}^{B},\\ \pi_{B}(R)(1-F(\tilde{i}(R)))+(1-\pi_{B}(R))F(\tilde{i}(R))&\text{if }\Delta_{S}^{B}<-2\Delta_{O}^{B},\end{cases}$ (8)
where we have made explicit that the indifferent receiver $\tilde{i}$ and the probability of a buy recommendation $\pi_{B}$ are functions of the threshold $R$. As for the value above, the relation between $\Delta_{S}^{B}$ and $\Delta_{O}^{B}$ determines which types accept a recommendation. In the first case, all types accept a recommendation so that demand is given by the probability of a buy recommendation $\pi_{B}(R)$. In the other cases, demand comprises both types that accept a recommendation and types that do not. Importantly, types that do not accept recommendations buy the recommended product following a don’t-buy recommendation, thus contributing to demand.
4 Symmetric Populations
In this section, we consider populations that are symmetrically distributed. Hence, there is no systematic preference bias for either of the controversial products in the population, so the preference heterogeneity reflects pure noise. We will first consider value-maximizing recommendation systems. Then, we turn to demand-maximizing recommendation systems. Throughout this section, we invoke the following symmetry assumption on the preference distribution.